I can not figure this out, thank you for your time
4. I am planning a party for my son’s birthday. I am planning on purchasing cookies for the children, along with brownies. Brownies come in packs of 6 and cookies come in packs of 18. Also, I can purchase a combination pack of 6 cookies and 6 brownies.
A. Write an expression to show how many cookies (c) and brownies (b) I purchased if I bought one pack of each.
B. Each cookie costs $.79 and each brownie costs $.59. How much would I pay for my three packs?

Answer by richwmiller(17219)

b6
c18
6b6c
3 packs
6b+18c+6b+6c=t
12b+24c=t
12*.59+24*.79=26.04